OpenStudy (anonymous):

A street magician asks a passerby to draw one card from a standard deck, remember it, and then replace it. The magician then shuffles the deck and asks the person to draw a second card. Find the probability that at least one of the two cards was a heart.

OpenStudy (dumbcow):

probability that neither card was a heart = 3/4 *3/4 = 9/16 prob at least 1 was heart = 1 - 9/16